Two angles are supplementary if the sum of their measures is 180 degrees the measure of the first angle is 12 degrees less than two times the second angle
what are the two angles

Answer by benazir.sj@gmail.com(109)

let the second angle be x
first angle is 12 degrees less than two times the second angle
=2x-12
given sum of there measures is 180
x+2x-12 =180
3x-12=180
3x=180+12
3x=192
x =64
second angle is 64
first angle will be
2x-12
2*64-12
128-12 =116
116 and 64 are the measures of the two angles